OpenStudy (anonymous):

E(n) is an even-number function. Find E(8) when E(n) = 2n

OpenStudy (anonymous):

HELLOW?????

OpenStudy (anonymous):

no w lol

Parth (parthkohli):

\[\textbf{HINT}\]\(\textit E (8) = 2(8)\)

Parth (parthkohli):

In better words, \(\textit E(8) = 2 \times 8\)

OpenStudy (anonymous):

16 yes yes but how do i work it out ???

OpenStudy (anonymous):

like how do i do this lol no teacher to help

OpenStudy (anonymous):

E(n) = E(8) so n = 8. Plug 8 into n so 2n = 2(8) = 16.

OpenStudy (anonymous):

I THINK I UNDERSTAND NOW !!!!!!!